A man cycles to his destination with the wind speed at 8m/s and returns home against the wind at 4m/s.What is his average speed?
1
Expert's answer
2019-09-02T09:26:39-0400
"{\rm Average\:speed}=\frac{{\rm Total\: distance}}{{\rm Total\: time}}"

"v=\frac{d_1+d_2}{t_1+t_2}=\frac{d+d}{d/v_1+d/v_2}=\frac{2v_1v_2}{v_1+v_2}"

"=\frac{2\times 8\times 4}{8+4}=5.3\:{\rm m/s}"